Kindly help me with this question.
The student-teacher ratio for Washington high was reported to be 27.5 to 1.
If there are 42 teachers, then how many students are there?
Thank you for your help.

Answer by stanbon(75887)

The student-teacher ratio for Washington high was reported to be 27.5 to 1.
If there are 42 teachers, then how many students are there?
x/42 = 27.5/1
x = 42*27.5 = 1155
=================================
Cheers,
Stan H.